OpenStudy (anonymous):

2. x²-6x+1=0 would the answer to this be x=-3+√10 AND x=-3-√10

jimthompson5910 (jim_thompson5910):

that is not correct

OpenStudy (anonymous):

can you help, I have no idea how to do it

jimthompson5910 (jim_thompson5910):

x^2 - 6x + 1 = 0 is in the form ax^2 + bx + c = 0 where a = 1 b = -6 c = 1

jimthompson5910 (jim_thompson5910):

see how I'm identifying a,b,c?

OpenStudy (anonymous):

yeah, I already did that

jimthompson5910 (jim_thompson5910):

now you plug those values into the quadratic formula \[\Large x = \frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

jimthompson5910 (jim_thompson5910):

So we get this \[\Large x = \frac{-b \pm \sqrt{b^2-4ac}}{2a}\] \[\Large x = \frac{-(-6) \pm \sqrt{(-6)^2-4(1)(1)}}{2(1)}\] What's next?

OpenStudy (anonymous):

got that, I got to -(-6)+-√36-(4) ---------------------------- 2

jimthompson5910 (jim_thompson5910):

then you'll have \[\Large x = \frac{6 \pm \sqrt{32}}{2}\]

jimthompson5910 (jim_thompson5910):

now you have to simplify \[\Large \sqrt{32}\]

OpenStudy (anonymous):

5.6?

jimthompson5910 (jim_thompson5910):

You simplify like this \[\Large \sqrt{32} = \sqrt{16*2}\] \[\Large \sqrt{32} = \sqrt{16}*\sqrt{2}\] \[\Large \sqrt{32} = 4\sqrt{2}\]

OpenStudy (anonymous):

so... x=4\[\sqrt{2}\]

jimthompson5910 (jim_thompson5910):

not quite

jimthompson5910 (jim_thompson5910):

\[\Large x = \frac{6 \pm \sqrt{32}}{2}\] will turn into \[\Large x = \frac{6 \pm 4\sqrt{2}}{2}\]

OpenStudy (anonymous):

x=10√2 ?

jimthompson5910 (jim_thompson5910):

you cannot add the 6 and 4 (since the 4 has a root 2 attached to it)

jimthompson5910 (jim_thompson5910):

what you can do though is factor out 2 from the numerator, then divide like this \[\Large x = \frac{6 \pm 4\sqrt{2}}{2}\] \[\Large x = \frac{2(3 \pm 2\sqrt{2})}{2}\] \[\Large x = 3 \pm 2\sqrt{2}\]

OpenStudy (anonymous):

so, x=3+√2 and x=3-√2?

jimthompson5910 (jim_thompson5910):

I would replace the 'and' with 'or' since x cannot be two values at once, but yes, the solutions are x=3+√2 or x=3-√2

OpenStudy (anonymous):

thanks so much!

jimthompson5910 (jim_thompson5910):

you're welcome